OpenStudy (anonymous):

What are the solutions of 6x2 − x − 15 = 0? A. x = 5/6, x = −3 B. x = −5/6, x = 3 C. x = 3/2, x = −5/6 D. x = −3/2, x = 5/6

OpenStudy (btaylor):

factor it into: (3x-5)(2x+3) = 0 now, set each factor equal to 0: 3x-5 = 0 --> x = 5/3 2x+3 = 0 --> x = -3/2 hmm...I factored right...I'd put D.

OpenStudy (anonymous):

Ok thank!